OpenStudy (anonymous):

i cant figure out the result of this integral using two methods: int ((a^2)+(x^2))^(-3/2)dx i used one method of substituting x with a*tan(theta). can anyone figure out another method? simpler is better. thanks

OpenStudy (lochana):

substitute x = a tan(theta)

OpenStudy (anonymous):

thankyou, this is what i have do but exists another method?

OpenStudy (lochana):

have you found the solution?

OpenStudy (anonymous):

yes the result is \[\frac{ x }{ a^2\sqrt{a^2+x^2} } +c\] but I want to use a method to compute the integral without using trigonometry

OpenStudy (anonymous):

if I use integration by parts how can I split the fuction:\[\int\limits_{}^{}f * g' dx\]